# SAM 2: Segment Anything in Images and Videos

### v1

@@ 6.2 Image tasks
The bottom row shows how training on our SA-1B and video data mix can further improve accuracy to 61.4% on average on the 23 datasets. We also see exceptional gains on the video benchmarks from SA-23 (video datasets are evaluated as images, identical to Kirillov et al. (2023)), and the 14 new video datasets we added.
@@
SAM 2 our mix 61.4 (83.7) 63.1 (83.9) 59.1 (83.3) 69.6 (86.0) 130.1
Table 6: Zero-shot accuracy on the Segment Anything (SA) task across 37 datasets. The table shows the average 1- and 5-click mIoU of SAM 2 compared to SAM by domains (image/video). We report the average metrics on the 23 datasets used by SAM (SA-23) and the average across 14 additional zero-shot video datasets (as detailed in §E.3).
Overall, the findings underscore SAM 2’s dual capability in interactive video and image segmentation, a strength derived from our diverse training data that encompasses videos and static images across visual domains. More detailed results including a breakdown by dataset are in §E.3.
@@ 7 Comparison to state-of-the-art in semi-supervised VOS
We also evaluate existing work on the SA-V val and test sets which measure performance for open-world segments of “any” object class. When comparing on this benchmark, we see that most previous methods peak at around the same accuracy. The best performance on SA-V val and SA-V test for prior work is significantly lower demonstrating the gap to a “segment anything in videos” capability. Finally, we see that SAM 2 also brings notable gains in long-term video object segmentation as observed in the LVOS benchmark result.
@@
8 Data and model ablations
This section presents ablations that informed the design decisions for SAM 2. We evaluate on our MOSE development set (“MOSE dev”) which contains 200 randomly-sampled videos from the MOSE training split and excluded from the training data in our ablations, SA-V val, and the average over 9 zero-shot video datasets. As the metric for comparison, we report $\mathcal{J}\&\mathcal{F}$ under 3-click input on the first frame as a balance between the 1-click regime and the VOS-style mask prompts. Additionally, we report the average 1-click mIoU on the 23-dataset benchmark used by SAM for the SA task on images. Unless otherwise specified, we run our ablations at 512 resolution and with SA-V manual and a 10% subset of SA-1B. Additional details are in §C.2.
8.1 Data ablations
Data mix ablation.
In Table 8, we compare the accuracy of SAM-2 when trained on different data mixtures. We pre-train on SA-1B and then train a separate model for each setting. We fix the number of iterations (200k) and batch size (128) with only the training data changing between experiments. We report accuracy on our SA-V val set, MOSE, 9 zero-shot video benchmarks, and the SA-23 tasks (§6.2). Row 1 shows that a model purely trained on VOS datasets (Davis, MOSE, YouTubeVOS) performs well on the in-domain MOSE dev, but poorly on all the others including the 9 zero-shot VOS datasets (59.7 $\mathcal{J}\&\mathcal{F}$).
We observe tremendous benefit from adding our data engine data into the training mix, including +12.1% average performance improvement on 9 zero-shot datasets (row 11 vs 1). This can be attributed to the limited coverage and size of VOS datasets. Adding SA-1B images improves the performance on the image segmentation task (rows 3 vs 4, 5 vs 6, 9 vs 10, 11 vs 12) without degrading the VOS capability. Training only on SA-V and SA-1B (row 4) is enough to obtain strong performance on all benchmarks except for MOSE. Overall, we obtain the best results when mixing all datasets: VOS, SA-1B, and our data engine data (row 12).
@@ 8.1 Data ablations
Figure 7: Performance of SAM 2 as a function of the SA-V quantity. We report $\mathcal{J}\&\mathcal{F}$ accuracy for 3-click prompts in the first frame on SA-V val (left), 9 zero-shot datasets (center), and MOSE dev (right).
@@
In Table 9, we experiment with filtering strategies for quality. We subsample 50k masklets from SA-V, either randomly or by taking the masklets that have been edited the most by annotators. Filtering based on the number of edited frames leads to strong performance using just 25% of the data, and outperforms random sampling. However, it is worse than using all 190k SA-V masklets.
@@
Setting SA-V val Intern-test MOSE dev 9 zero-shot SA-23
@@
8.2 Model architecture ablations
@@ 8.2.1 Capacity ablations
Table 10: Capacity ablations. We ablate modeling capacity along input size (resolution, #frames), memory size (#memories, memory channel dim) and model size (memory attention, image encoder). Ablation defaults in gray.
@@
During training, we sample sequences of frames of fixed resolution and fixed length (here denoted by # frames). We ablate their impact in Tables LABEL:tab:tab-results-input-resolution, LABEL:tab:tab-results-input-frames. A higher resolution leads to significant improvements across image and video tasks, and we use an input resolution of 1024 in our final model. Increasing the number of frames brings notable gains on video benchmarks and we use a default of 8 to balance speed and accuracy.
@@
More capacity in the image encoder or memory-attention (#self-/#cross-attention blocks) generally leads to improved results, as shown in Tables LABEL:tab:tab-results-memory-attention-size, LABEL:tab:tab-results-image-encoder. Scaling the image encoder brings gains on both image and video metrics, while scaling the memory-attention only improves video metrics. We default to using a B+ image encoder, which provides a reasonable balance between speed and accuracy.
8.2.2 Relative positional encoding
@@
RPB in img. enc. 2d-RoPE in mem. attn. MOSE dev SA-V val LVOSv2 val 9 zero-shot speed SA-23
@@
Table 11: Relative positional encoding. We use 2d-RoPE in memory attention while removing RPB from the image encoder by default (gray). Removing RPB also allows us to enable FlashAttention-2 (Dao, 2023), which gives a significant speed boost at 1024 resolution. At the higher resolution of 1024, the speed gap between 2d-RoPE (1st row) and the no RoPE baseline (3rd row) becomes much smaller.
@@
8.2.3 Memory architecture ablations
@@
We investigate the effectiveness of feeding the memory features to a GRU before adding them to the memory bank. Similar to §8.2.2, we also evaluate on LVOSv2 as an additional benchmark for long-term object segmentation. While prior works have commonly employed GRU (Cho et al., 2014) states as a means of incorporating memory into the tracking process, our findings in Table 12 suggest that this approach does not provide an improvement (except slightly on LVOSv2). Instead, we find it sufficient to directly store the memory features in the memory bank, which is both simpler and more efficient.
@@
We ablate the impact of cross-attending to the object pointer vectors from the mask decoder output in other frames (see §4). The results presented in Table 12 show that while cross-attending to object pointers does not enhance average performance across the 9 zero-shot datasets, it significantly boosts performance on SA-V val dataset as well as on the challenging LVOSv2 benchmark (validation split). Hence, we default to cross-attending to object pointers together with the memory bank.
@@
9 Conclusion
We present a natural evolution of Segment Anything into the video domain, based on three key aspects: (i) extending the promptable segmentation task to video, (ii) equipping the SAM architecture to use memory when applied to video, and (iii) the diverse SA-V dataset for training and benchmarking video segmentation. We believe SAM 2 marks a significant advancement in visual perception, positioning our contributions as milestones that will propel further research and applications in the field.
10 Acknowledgements
Appendix
Table of contents:
  • §A: Task Details
  • §B: Limitations
  • §C: Model Details
  • §D: Dataset Details
  • §E: Zero-shot Experiments Details
  • §G: Dataset, Annotation, and Model Cards
  • §D.2.1: Annotation Guidelines
Appendix A Details on the PVS Task
@@
PVS is related to several tasks in both the static image and video domains. On images, the SA task can be considered a subset of PVS with the video reduced to a single frame. Similarly, traditional semi-supervised and interactive VOS (Pont-Tuset et al., 2017) tasks are special cases of PVS, limited to mask prompts provided only on the first frame and scribbles on multiple frames to segment objects throughout a video, respectively. In PVS, prompts can either be clicks, masks, or boxes, and the focus is on enhancing the interactive experience, enabling easy refinement of an object’s segmentation with minimal interaction.
Appendix B Limitations
@@
Appendix C SAM 2 details
C.1 Architecture
@@
The prompt encoder design follows SAM, and we next discuss additional details on design changes in the mask decoder. We use the mask token corresponding to the output mask as the object pointer token for the frame, which is placed in the memory bank. As discussed in §4, we also introduce an occlusion prediction head. This is accomplished by including an additional token along with the mask and IoU output tokens. An additional MLP head is applied to this new token to produce a score indicating the likelihood of the object of interest being visible in the current frame (as shown in Figure 9).
@@
C.2 Training
C.2.1 Pre-training
We first pre-train SAM 2 on static images on the SA-1B dataset (Kirillov et al., 2023). Table LABEL:tab:pre-training-hparam details the settings used during pre-training on SA-1B – other settings not mentioned here follow Kirillov et al. (2023). The image encoder is initialized from MAE pre-trained Hiera (Ryali et al., 2023). Similar to SAM, we filter masks covering more than 90% of the image and restricted training to 64 randomly sampled masks per image.
@@ C.2.2 Full training
SAM 2 is designed for two tasks; the PVS task (on videos) and the SA task (on images). Training is done jointly on image and video data. To optimize our data usage and computational resources during training, we adopt an alternating training strategy between video data (multiple frames) and static images (one single frame). Specifically, in each training iteration, we sample a full batch either from the image or video dataset, with their sampling probabilities proportional to the size of each data source. This approach allows for a balanced exposure to both tasks and a different batch size for each data source to maximize compute utilization. Settings not explicitly mentioned here for the image task follow settings from the pre-training phase. See Table 13 (b) for the hyperparameters in our full training stage. The training data mixture consists of $\sim$15.2% SA-1B, $\sim$70% SA-V and $\sim$14.8% Internal. The same settings are used when open-source datasets are included, with the change that the additional data is included ($\sim$1.3% DAVIS, $\sim$9.4% MOSE, $\sim$9.2% YouTubeVOS, $\sim$15.5% SA-1B, $\sim$49.5% SA-V, $\sim$15.1% Internal).
@@
warmup linear, 15k iters
@@
img. augmentation hflip, resize to 1024 (square)
vid. aug. hflip, affine (deg: 25, shear: 20), colorjitter (b: 0.1, c: 0.03, s: 0.03, h: null), grayscale (0.05), per frame colorjitter (b: 0.1, c: 0.05, s: 0.05, h: null)
@@
We restrict the maximum number of masklets for each sequence of 8 frames to 3 randomly chosen ones. We reverse the temporal order with a probability of 50% to help generalization to bi-directional propagation. When we sample corrective clicks - with a small probability of 10%, we randomly sample clicks from the ground truth mask, irrespective of the model prediction, to allow additional flexibility in mask refinement.
@@ C.3 Speed benchmarking
Appendix D Data details
D.1 SA-V dataset details
@@
D.2 Data engine details
D.2.1 Annotation protocol

### v2

@@ 6.3 Image segmentation
The bottom row shows how training on our SA-1B and video data mix can further improve accuracy to 61.4% on average on the 23 datasets. We also see exceptional gains on the video benchmarks from SA-23 (video datasets are evaluated as images, identical to Kirillov et al. (2023)), and the 14 new video datasets we added. More detailed results including a breakdown by dataset are in §F.3.
@@
SAM 2 our mix 61.9 (83.5) 63.3 (83.8) 60.1 (83.2) 69.6 (85.8) 130.1
Table 5: Zero-shot accuracy on the Segment Anything (SA) task across 37 datasets. The table shows the average 1- and 5-click mIoU of SAM 2 compared to SAM by domains (image/video). We report the average metrics on the 23 datasets used by SAM (SA-23) and the average across 14 additional zero-shot video datasets (as detailed in §F.3).
@@ 7 Comparison to state-of-the-art in semi-supervised VOS
We also evaluate existing work on the SA-V val and test sets which measure performance for open-world segments of “any” object class. When comparing on this benchmark, we see that most previous methods peak at around the same accuracy. The best performance on SA-V val and SA-V test for prior work is significantly lower demonstrating the gap to a “segment anything in videos” capability. Finally, we see that SAM 2 also brings notable gains in long-term video object segmentation as observed in the LVOS benchmark result. For data and model ablations, see §A.
8 Conclusion
We present a natural evolution of Segment Anything into the video domain, based on three key aspects: (i) extending the promptable segmentation task to video, (ii) equipping the SAM architecture to use memory when applied to video, and (iii) the diverse SA-V dataset for training and benchmarking video segmentation. We believe SAM 2 marks a significant advancement in visual perception, positioning our contributions as milestones that will propel further research and applications.
Acknowledgements.
Appendix
Table of content:
  • §A: Data and Model Ablations
  • §B: Task Details
  • §C: Limitations
  • §D: Model Details
  • §E: Dataset Details
  • §E.2.1: Annotation Guidelines
  • §F: Zero-shot Experiments Details
  • §H: Dataset, Annotation, and Model Cards
Appendix A Data and model ablations
This section presents ablations that informed the design decisions for SAM 2. We evaluate on SA-V val, Internal-test, our MOSE development set (“MOSE dev”) which contains 200 randomly-sampled videos from the MOSE training split, excluded from our training data and the average over 9 zero-shot video datasets. As the metric for comparison, we report $\mathcal{J}\&\mathcal{F}$ under 3-click input on the first frame as a balance between the 1-click regime and the VOS-style mask prompts. Additionally, we report the average 1-click mIoU on the 23-dataset benchmark used by SAM for the SA task on images. Unless otherwise specified, we perform our ablations at 5122 spatial resolution, trained with SA-V manual and a 10% subset of SA-1B. Additional details are in §D.2.
A.1 Data ablations
@@
Data mix ablation.
In Table 7, we compare the accuracy of SAM 2 when trained on different data mixtures. We pre-train on SA-1B and then train a separate model for each setting. We fix the number of iterations (200k) and batch size (128) with only the training data changing between experiments. We report accuracy on our SA-V val and Internal set, MOSE, 9 zero-shot video benchmarks, and the SA-23 tasks (§6.3).
Row 1 shows that a model purely trained on existing VOS datasets (DAVIS, MOSE, YouTubeVOS) performs well on the in-domain MOSE dev, but poorly on all the others including the 9 zero-shot VOS datasets (59.7 $\mathcal{J}\&\mathcal{F}$). We observe tremendous benefit from adding our data engine data into the training mix, including +12.1% average performance improvement on 9 zero-shot datasets (row 11 vs 1). This can be attributed to the limited coverage and size of VOS datasets. Adding SA-1B images improves the performance on the image segmentation task (rows 3 vs 4, 5 vs 6, 9 vs 10, 11 vs 12) without degrading the VOS capability. Training only on SA-V and SA-1B (row 4) is enough to obtain strong performance on all benchmarks except for MOSE (specific object categories). Overall, we obtain the best results when mixing all datasets: VOS, SA-1B, and our data engine data (row 12).
@@
Figure 6: SAM 2 accuracy as a function of the SA-V quantity. We report $\mathcal{J}\&\mathcal{F}$ accuracy for 3-click prompts in the first frame on SA-V val (left), 9 zero-shot datasets (center), and MOSE dev (right).
@@
In Table 8, we experiment with filtering strategies for quality. We subsample 50k masklets from SA-V, either randomly or by taking the masklets that have been edited the most by annotators. Filtering based on the number of edited frames leads to strong performance using just 25% of the data, and outperforms random sampling, but is worse than using all 190k SA-V masklets.
@@
Setting SA-V val Internal-test MOSE dev 9 zero-shot SA-23
@@
A.2 Model architecture ablations
@@ A.2.1 Capacity ablations
Table 9: We ablate modeling capacity along input size (resolution, #frames), memory size (#memories, memory channel dim) and model size (memory attention, image encoder). Ablation defaults in gray.
@@
During training, we sample sequences of frames of fixed resolution and fixed length (here denoted by # frames). We ablate their impact in Tables LABEL:tab:tab-results-input-resolution, LABEL:tab:tab-results-input-frames. A higher resolution leads to significant improvements across image and video tasks, and we use a spatial input resolution of 10242 in our final model. Increasing the number of frames brings notable gains on video benchmarks and we use a default of 8 to balance speed and accuracy.
@@
More capacity in the image encoder or memory-attention (#self-/#cross-attention blocks) generally leads to improved results, as shown in Tables LABEL:tab:tab-results-memory-attention-size, LABEL:tab:tab-results-image-encoder. Scaling the image encoder brings gains on both image and video metrics, while scaling the memory-attention only improves video metrics. We default to using a B+ image encoder, which provides a reasonable balance for speed and accuracy.
A.2.2 Relative positional encoding
@@
RPB 2d-RoPE MOSE dev SA-V val LVOSv2 val 9 zero-shot speed SA-23
@@
Table 10: We use 2d-RoPE positional encoding in memory attention while removing RPB from the image encoder by default (gray). Removing RPB also allows us to enable FlashAttention-2 (Dao, 2023), which gives a significant speed boost at (10242) resolution. At such higher resolution , the speed gap between 2d-RoPE (1st row) and the no RoPE baseline (3rd row) becomes small (4%).
@@
A.2.3 Memory architecture ablations
@@
We investigate the effectiveness of feeding the memory features to a GRU before adding them to the memory bank. Similar to §A.2.2, we also evaluate on LVOSv2 as an additional benchmark for long-term object segmentation. While prior works have commonly employed GRU (Cho et al., 2014) states as a means of incorporating memory into the tracking process, our findings in Table 11 suggest that this approach does not provide an improvement (except slightly on LVOSv2). Instead, we find it sufficient to directly store the memory features in the memory bank, which is both simpler and more efficient.
@@
We ablate the impact of cross-attending to the object pointer vectors from the mask decoder output in other frames (see §4). The results presented in Table 11 show that while cross-attending to object pointers does not enhance average performance across the 9 zero-shot datasets, it significantly boosts performance on SA-V val dataset as well as on the challenging LVOSv2 benchmark (validation split). Hence, we default to cross-attending to object pointers together with the memory bank embeddings from the memory encoder.
@@
Appendix B Details on the PVS Task
@@
PVS is related to tasks in the image and video domains. For images, the SA task can be considered a subset of PVS with the video reduced to a single frame. Similarly, traditional semi-supervised and interactive VOS (Pont-Tuset et al., 2017) tasks are special cases of PVS, limited to mask prompts provided only on the first frame and scribbles on multiple frames to segment objects throughout a video, respectively. In PVS, prompts can either be clicks, masks, or boxes, and the focus is on enhancing the interactive experience, enabling refinement of a segmentation with minimal interaction.
Appendix C Limitations
@@
Appendix D SAM 2 details
D.1 Architecture
@@
The prompt encoder design follows SAM, and we next discuss additional details on design changes in the mask decoder. We use the mask token corresponding to the output mask as the object pointer token for the frame, which is placed in the memory bank. As discussed in §4, we also introduce an occlusion prediction head. This is accomplished by including an additional token along with the mask and IoU output tokens. An additional MLP head is applied to this new token to produce a score indicating the likelihood of the object of interest being visible in the current frame (as shown in Figure 8). In the memory bank, we also add a learned occlusion embedding to the memory features of those frames that are predicted to be occluded (invisible) by the occlusion prediction head.
@@
Handling multiple objects in a video.
When applying SAM 2 to segment multiple objects in the same video (such as multi-object tracking in the semi-supervised VOS evaluation), we perform inference on each object independently. More specifically, we share the visual features from the image encoder between all the objects in the video but run all the other model components (such as the memory bank and the mask decoder) separately for each object.
D.2 Training
D.2.1 Pre-training
We first pre-train SAM 2 on static images on the SA-1B dataset. Table LABEL:tab:pre-training-hparam details the settings used during pre-training on SA-1B – other settings not mentioned here follow Kirillov et al. (2023). The image encoder is initialized from MAE pre-trained Hiera (Ryali et al., 2023). Similar to SAM, we filter masks covering more than 90% of the image and restricted training to 64 randomly sampled masks per image.
@@ D.2.2 Full training
SAM 2 is designed for two tasks; the PVS task (on videos) and the SA task (on images). Training is done jointly on image and video data. To optimize our data usage and computational resources during training, we adopt an alternating training strategy between video data (multiple frames) and static images (one single frame). Specifically, in each training iteration, we sample a full batch either from the image or video dataset, with their sampling probabilities proportional to the size of each data source. This approach allows for a balanced exposure to both tasks and a different batch size for each data source to maximize compute utilization. Settings not explicitly mentioned here for the image task follow settings from the pre-training phase. See Table 12 (b) for the hyperparameters in our full training stage. The training data mixture consists of $\sim$15.2% SA-1B, $\sim$70% SA-V and $\sim$14.8% Internal. The same settings are used when open-source datasets are included, with the change that the additional data is included ($\sim$1.3% DAVIS, $\sim$9.4% MOSE, $\sim$9.2% YouTubeVOS, $\sim$15.5% SA-1B, $\sim$49.5% SA-V, $\sim$15.1% Internal). When training on SA-V and other video datasets, we only use those manually annotated masklets (without adding automatically generated ones), which are sufficient to achieve strong performance based on our analyses.
We apply a series of data augmentations to the training videos (detailed in Table 12), including random horizontal flips, random affine transforms, random color jittering, and random grayscale transforms, as listed in Table 12. We also adopt a mosaic transform to simulate challenging scenarios with multiple similar-looking objects – with 10% probability, we tile the same training video into a $\times2$ grid and select a masklet from one of the 4 quadrants as the target object to segment. In this case, the model must focus on other cues like motion or temporal continuity to distinguish the target object from their identical-looking counterparts in other quadrants. In addition, the videos and objects in each quadrant are smaller in size (only half the original width and height) after this mosaic transform, which also facilitates learning to segment small objects.
@@
warmup linear, 7.5k iters
@@
image augmentation hflip, resize to 1024 (square)
video augmentation hflip, affine (deg: 25, shear: 20), colorjitter (b: 0.1, c: 0.03, s: 0.03, h: null), grayscale (0.05), per frame colorjitter (b: 0.1, c: 0.05, s: 0.05, h: null), mosaic-$\times2$ (0.1)
@@
We restrict the maximum number of masklets for each sequence of 8 frames to 3 randomly chosen ones. We reverse the temporal order with a probability of 50% to help generalization to bi-directional propagation. When we sample corrective clicks, with a small probability of 10%, we randomly sample clicks from the ground truth mask, irrespective of the model prediction, to allow additional flexibility in mask refinement.
Fine-tuning using 16-frame sequences.
A potential shortcoming of the procedure above is that the model only sees sampled 8-frame sequences during training, which is relatively short compared to the full video length during inference. To alleviate this issue and further boost the segmentation quality on long videos, we introduce an extra fine-tuning stage where we sample 16-frame sequences on challenging videos (those videos with the highest number of edited frames, as described in §E.2.1). More specifically, we sort our masklets by number of edited frames and only consider the top 50% most edited masklets for training, for both SA-V and Internal datasets. We still keep the complete versions of the OSS datasets (DAVIS, MOSE, and YouTubeVOS) in the training mix. We fine-tune for 50k iterations (1/3 of the original schedule) using half of the original learning rate and freeze the image encoder to fit the 16-frame sequence into the 80 GB memory of A100 GPUs.
@@ D.3 Speed benchmarking
Appendix E Data details
E.1 SA-V dataset details
@@
Dataset diversity. As shown in Fig. 10, SA-V videos were recorded across 47 countries (Fig. 10(b)), by diverse participants (self-reported demographics in Fig. 10(c)). Fig. 10(a) shows a comparison of mask size distribution (normalized by video resolution) with DAVIS, MOSE, and YouTubeVOS. More than $88\%$ of SA-V masks have a normalized mask area less than 0.1.
@@
(a) Size
(b) Geography
Male274Female236
i Gender
18-2410925-4030541-6488
$i$ Age
(c) Crowdworker Demographics
E.1.1 Fairness evaluation
1-click 3-click mask
gender
male 81.9 95.1 95.9
female 75.1 94.1 95.2
age
18-26 77.2 95.0 95.7
26-50 76.7 94.7 95.8
50+ 81.4 95.1 96.2
In Appendix H, we provide model, data and annotation cards for SA-V.
E.2 Data engine details
E.2.1 Annotation protocol